OpenStudy (anonymous):

Find the value of x and y. A) x=81, y=98 B) x=71, y=98 C) x=81, y=108 D) x=71, y=108

OpenStudy (anonymous):

Is there an image?

OpenStudy (anonymous):

Well because they are supplementary, x has to equal 71

OpenStudy (anonymous):

and in all quadrilaterals, the interior angles add up to 360, so y equals?

OpenStudy (anonymous):

100+81+71=252

OpenStudy (anonymous):

360-252= 108

OpenStudy (anonymous):

y=108

OpenStudy (anonymous):

i thought it was 98 at first but now i see how you got 108 lol

OpenStudy (anonymous):

Anymore questions?

OpenStudy (anonymous):

yes lol hold on a sec tho .

OpenStudy (anonymous):

Find the value of x. The diagram is not drawn to scale. A)14 B)15 C)16 D) 17

OpenStudy (anonymous):

I want you to try solving this, after the few we did together

OpenStudy (anonymous):

hint: start with 180(n-2)

OpenStudy (anonymous):

okay , so would i replace n with how many sides there are on this shape ?

OpenStudy (anonymous):

Yep, so 5

OpenStudy (anonymous):

so , 180(5-20)= 540. what do I do after that.. lol

OpenStudy (anonymous):

I meant 2 not 20.

OpenStudy (anonymous):

Sorry internet cut out

OpenStudy (anonymous):

Ok so you know already about the first 340 degrees of the pentagon

OpenStudy (anonymous):

so 540-340=200

OpenStudy (anonymous):

so what's left is 7x+8 and 5x

OpenStudy (anonymous):

add them to get 12x+8

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

kinda lol it's all really confusing to meee.

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

i don't understand the part right before you said are you following ? i don't understand how you added or what you did to get 12x+8

OpenStudy (anonymous):

You know about the first 340 degrees

OpenStudy (anonymous):

you are missing 200

OpenStudy (anonymous):

the measurements you don't know are 7x+8 and 5x

OpenStudy (anonymous):

but you know that 200 is 7x+8 and 5x added

OpenStudy (anonymous):

so when you add you get.... 12x+8

OpenStudy (anonymous):

oh , okay i understand how you got it . lol

OpenStudy (anonymous):

so now subtract 8 from both sides, you get 12x=192 And Finnish off by dividing both sides by 12 and getting?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

I'm getting off, I hope I helped

OpenStudy (anonymous):

lol you did , thank you :)